Electrons are emitted with zero velocity from metal surface when it is exposed to radiation of wavelength 7000 angstrom. Calculate 3 sholed frequency,work function of the metal?

Grade:11

2 Answers

Eshan
5 years ago
Dear student,

Since the electrons are emitted with no kinetic energy, all the energy must have been spent in the work function of the metal.

HenceW_0=h\nu=h\dfrac{c}{\lambda}=1.77eV
Arun
5 years ago
c = νλ∴ νo = c/λo = 3.0×10^8 ms-1/7000×10^-10 m = 4.28×10^14 s-1Work function (wo) = hνo = 6.626×10^-34 Js×4.28×10^14 s-1 = 2.835×10^-19 J